OpenStudy (korosh23):

What is the inverse of the function f(x) = (x+2) ^2 ???

OpenStudy (korosh23):

I believe the answer is: \[f^-1 (x)= \sqrt{x -2}\]

OpenStudy (korosh23):

Am I right?

OpenStudy (astrophysics):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (astrophysics):

\[y=(x+2)^2 \implies x = (y+2)^2\] \[y+2=\sqrt{x} \implies y = \sqrt{x}-2\]

OpenStudy (astrophysics):

I think you may have made a mistake sqaure rooting both sides

OpenStudy (astrophysics):

\[f^{-1}(x)=\sqrt{x}-2\]

OpenStudy (astrophysics):

I guess we should put \[f^{-1}(x) = \pm \sqrt{x}-2\] if you really want to be precise

OpenStudy (korosh23):

why the square root is not part of the -2 ? @Astrophysics

OpenStudy (rajat97):

look we just need to find x in terms of y so we have y=(x+2)^2 so sqrt(y) = x+2 then, sqrt(y) -2 = x now just interchange x and y so, you'll get, sqrt(x) - 2 = y it seems to be like this:\[y=\sqrt{x} -2\]

OpenStudy (korosh23):

ok I got it thank you

OpenStudy (rajat97):

you're welcome